OpenStudy (anonymous):

Use the distributive property to simplify the expression below: -2(2u - 1)

OpenStudy (anonymous):

-4u+2

OpenStudy (anonymous):

okay so you would start bu multiplying -2 to the 2u and the -1.

OpenStudy (anonymous):

The distributive property is: \[c(a+b) = ca +cb\] can you identify a, b, and c?

OpenStudy (anonymous):

no please explain

OpenStudy (anonymous):

i could explain...so

OpenStudy (anonymous):

If we put it into the ca + cb form, we get -2(2u) + -2(-1)

OpenStudy (anonymous):

In this case c = -2, a = 2u and b = -1. If we were to but it back into c(a+b), we get -2(2u+-1), this is the same as -2(2u-1)

OpenStudy (anonymous):

two minus give a plus, you should multiply a digit or symbol to each symbol or digit in the parentheses

OpenStudy (anonymous):

@wio thank you @vacuole you made it very clear thankx

OpenStudy (anonymous):

-2(2u-1)=-2*2u-(-2)*1=-4u+2